Problem 21.81

Consider an RC circuit with E = 12.0 V ,R = 195 Ω , and C = 55.8 μF .

Part A

Find the time constant for the circuit.

τ =   ms  

SubmitMy AnswersGive Up

Part B

Find the maximum charge on the capacitor.

qmax =   μC  

SubmitMy AnswersGive Up

Part C

Find the initial current in the circuit.

I(0) =   mA  

SubmitMy AnswersGive Up
Answer #1

A) Time constant, T = R*C

= 195*55.8*10^-6

= 1.088*10^-2 s

= 10.88 ms

B) Qmax = C*V

= 55.8*10^6*12

= 669.6 micro C

C) Io = V/R

= 12/195

= 61.5 mA